LHC 2010 - latest news

Week 47: Machine coordinators: Ralph Assmann– Jan Uythoven

Main aims of the week:

LHC coordination vistar



Link to report from morning meeting

Monday 29th November

  • 00h48: Start ramp.
  • 02h34: Stable beams #1528. 121b x 121b.
  • 03h45: Beams dumped. RB.A78 in slow abort. PO piquet and MP3 called. Problem related to the new feature that superlocked circuits?
  • 05h46: Injection.
  • 07h16: Start ramp. 121b x 121b.
  • 07h55: Beams dumped during squeeze. RB A67 is in slow power abort. In PIC application it is superlocked.

Slides from 8:30 meeting

Sunday 28th November

  • 00h00: Stable beams #1525 121 x 121 b.
    Peak lumi 2.0e25 - 2.3e25 cm-2 s-1.
  • 09h52: Dump fill #1525 - delivered 440 (mb)-1. RCBCH8.L5B1 tripped during the rampdown precycle.
  • 10h49: Injection.
  • 12h21: Start ramp.
  • 13h22: Stable beams #1526. 121b x 121b. Turnaround: 3h30. Length: 5h35.
  • 18h57: Beams dumped by request.
  • 19h56: Injection. BPM orbit excursion interlock could not be reset with a two low-intensity bunches in last batch -> dumped. Orbit feedback problem.

Slides from 9:00 meeting

Saturday 27th November

  • 00h10: Injection. Some time spent to sort out chromaticity.
  • 03h00: Start ramp.
  • 04h01: Stable beams #1523. 121b x 121b. Peak lumi: 2.8e25 - 3.2e25 cm-2 s-1. Length: 8h23.
  • 12:24 Dump fill #1523
  • 21:30 Injecting for fill #1525

Slides from 9:00 meeting

Friday 26th November

  • 00h00: Beams dumped, lost sector 81.
  • 03h30: Start refill #1521
  • 05h44: Stable beams #1521. 121b x 121b. Length: 4h10.
  • 09h54: Beam dump. Instability on electrical network (RD.LR1 tripped). ATLAS solenoid ramp down.
  • 11h01: Injection.
  • 12h28: Start ramp.
  • 13h32: Stable beams #1522. 121b x 121b. Turnaround: 3h38. Luminosity: 2.8e25 cm-2 s-1. Length: 8h09.
  • 21h41: Beams dumped on request. BPM calibration needed to correct temperature drift on the electronics in SR7. CV piquet has installed temperature monitoring at the crates concerned and will monitor the temperature over the weekend. At the moment the temperature is 22 degrees.
  • 22h18: ALICE access required to fix trigger problem on half of their muon chambers.

Slides from 8:30 meeting

Thursday 25th November

  • 03h34: Beams dumped on request. Length=5h31. Rampdown and precycle.
  • 04h20: Injection. Losses in SPS flat bottom reduced. Now losses at acceleration had increased and source went down.
  • 05h54: Start ramp.
  • 06h58: Stable beams #1518. 121b x 121b. Turn-around: 3h24.
  • 07h16: CMS sees cogging problem. Other experiments fine. Suspect threshold effect on BPTX signal (low bunch intensity).
  • 08h06: Beam dump: Q9.R2 à noise on long cables (seen that before).
  • Work on BGI, RF trim and injector chain
  • 16h15: Injection fill # 1520
  • 18h12: Stable beams #1520. 121b x 121b.

Slides from 8:30 meeting

Wednesday 24th November

  • 02h04: Stable beams #1514. 121b x 121b.
  • 08h32: Beams dumped on request. Length=6h28. Rampdown and precycle.
  • 09h41: Injection.
  • 12h40: Start ramp.
  • 14h01: Stable beams #1515. 121b x 121b. Turn-around: 5h29.
  • 17h01: Beams dumped on request. Length=3h00. Rampdown and precycle.
  • 18h04: Injection. Correction injection oscillations.
  • 18h40: Trip much of sector 81. Glitch on temperature sensor on RSS.A81.B1 (which is already superlock!).
  • 19h56: Injection.
  • 21h04: Start ramp.
  • 22h03: Stable beams #1517. 121b x 121b. Turn-around: 3h46 (1h16 subtracted for S8 trip).
  • 22h50: Lost QPS-OK on RCS.A45B1.

Slides from 8:30 meeting

Tuesday 23rd November

  • Change of ALICE polarity
  • Fill for stable beams overnight, fill #1514

Slides from 8:30 meeting

Monday 22nd November

  • 01:30 Stable beams #1506

Slides from 8:30 meeting